### v1

# Thermodynamics of Black Holes, far from Equilibrium

Abstract
As in thermodynamics, the celebrated first law of black hole mechanics relates infinitesimal changes in the properties of nearby equilibrium states of black holes (without reference to any physical process that causes the transition). The second law is a qualitative statement that the area of an event horizon cannot decrease under appropriate physical assumptions. These laws are generalized. The new first law applies to black holes in general relativity that can be arbitrarily far from equilibrium and refers to finite changes that occur due to physical processes. The new second law is a quantitative statement that relates the change in the dynamical horizon area with the flux of energy falling into the black hole in a physical process.
Introduction:   Some five decades ago, Bardeen, Carter and Hawking (BCH) [1, 2] discovered the first law for Killing horizons (KHs) that governs the relation between two nearby stationary, axisymmetric black holes (BHs) in general relativity. Hawking also showed that, in dynamical situations, the area of an event horizon (EH) cannot decrease if the space-time is asymptotically predictable and satisfies the null energy condition [3]. The close similarities with the first and second laws of thermodynamics, and Hawking’s subsequent discovery [4] that BHs emit quantum radiation at a temperature $(\kappa\hbar)/2\pi$ led to the identification of $A/4G\hbar$ as BH entropy. These discoveries have continued to inspire investigations on the quantum nature of BHs for half a century!
However, EHs are teleological; they can form and grow in flat regions of space-time (see, e.g., [5, 6, 7]). To determine if a given space-time admits an EH one needs to know the metric all the way to the infinite future. Therefore, their use in dynamical situations leads to serious difficulties, both at practical and conceptual levels. At the practical level, one cannot use EHs e.g. in the course of numerical simulations to locate the progenitor BHs, nor the remnant; they can only be introduced as an ‘afterthought’ once the simulation is complete. In the investigation of conceptual issues –such as cosmic censorship in classical general relativity [8], or the endpoint of the evaporation process of BHs in quantum gravity [9]– one cannot make a priori assumptions about the global nature of space-time. Hence, one does not know whether EHs even exist in interesting situations [10]. What, then, should BH entropy refer to in dynamical situations?
An answer is provided by quasi-local horizons (QLHs), introduced over two decades ago (see, in particular, [11, 12, 13, 14, 15, 16, 5, 17, 18, 19, 20, 21].) Since they are free of teleology, they are routinely used in numerical simulations of BH mergers. Their properties have been investigated using geometric analysis (see, e.g.,[22, 23, 24, 8, 25, 26]) as well as numerical methods [27, 28, 29, 30, 31], and they have also been used to analyze the BH evaporation process (see, e.g., [32, 33, 34, 35, 36, 37]). In this Letter, we will show that QLHs also allow for an extension of the first law to fully dynamical situations, as well as a more physical, quantitative version of the second law.
The First Law and its Reformulation:   In standard treatments, one begins with asymptotically flat space-times $(M,g_{ab})$ that admit a time translation Killing field $\mathring{t}^{a}$, a rotational Killing field $\mathring{\varphi}^{a}$, and a KH $H$ to which a (constant) linear combination $\mathring{k}^{a}:=\mathring{t}^{a}+\mathring{\Omega}\,\mathring{\varphi}^{a}$ of the two Killing fields serves as the null normal. The freedom in rescaling $\mathring{k}^{a}$ by a constant is fixed by requiring that $\mathring{t}^{a}$ be the unit time translation at infinity (and the affine parameter of $\mathring{\varphi}^{a}$ run from 0 to $2\pi$). Because of stationarity, of the ten Einstein’s equations only the four constraints are non-trivial. They imply that a number of identities must hold between geometrical quantities evaluated at $H$ and those evaluated at spatial infinity. Among them is the celebrated BCH first law: $\delta M=(\mathring{\kappa}/8\pi G)\delta A+\mathring{\Omega}\,\delta J$. Here $M,J$ are taken to be the Arnowitt-Deser-Misner (ADM) quantities, evaluated at infinity, while $A$ is the horizon area, and $\mathring{\kappa}$ the surface gravity of $\mathring{k}^{a}$, both evaluated at the horizon.
This formulation of the first law has an unsatisfactory feature [13, 14]: the normalization condition on $\mathring{t}^{a}$ and the ADM charges $M,J$ refer to space-time geometry near spatial infinity, far from the black hole. One would expect the laws of BH mechanics to refer just to the black hole. Indeed, quantities that enter the first law of thermodynamics refer only to the system under consideration. In dynamical situations, one has to consider physical processes that change the mass and angular momentum of the BH. Then the limitation becomes much more severe because the ADM $M$ and $J$ refer to the total system and are absolutely conserved in physical processes! $\delta M$ and $\delta J$ in the first law should now refer to infalling energy and angular momentum across the horizon of each BH in the given space-time, without reference to infinity.
However, thanks to the BH uniqueness theorem [38, 39], one can recast the standard first law using quantities that are specified just at $H$. First note that the horizon areal radius $R$ (defined by $A=4\pi R^{2}$), and the horizon angular momentum $J_{H}=-\frac{1}{8\pi G}\oint\epsilon_{abcd}(\nabla^{a}\varphi^{b})\,{\rm d}S^{cd}$ (the Komar integral of $\mathring{\varphi}^{a}$ at $H$) are determined by fields at the horizon. Secondly, a Kerr black hole can be characterized by values of $(R,J_{H})$. Next, since the surface gravity $\mathring{\kappa}$ scales linearly with the horizon null normal, one can fix the rescaling freedom in $\mathring{k}^{a}$ by requiring that its surface gravity $\mathring{\kappa}$ –also determined by the horizon geometry– be given by $\mathring{\kappa}=\kappa_{{}_{\rm Kerr}}(R,J)$. As noted above, there is no rescaling freedom in $\mathring{\varphi}^{a}$. Therefore by setting $\mathring{\Omega}=\mathring{\Omega}_{{}_{\rm Kerr}}(R,J)$, one can recover the restriction of $\mathring{t}^{a}$ to $H$ via $\mathring{t}^{a}\,\hat{=}\,\mathring{k}^{a}-\mathring{\Omega}\mathring{\varphi}^{a}$. (Throughout $\hat{=}\,$ will stand for equality at the QLH under consideration.) Then, the three (Komar) charges evaluated at the horizon are given by $Q_{{}^{(\mathring{k})}}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G},\,Q_{(\mathring{\varphi})}\,\hat{=}\,-2J_{H}$  and  $Q_{(\mathring{t})}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G}+2\mathring{\Omega}J_{H}$. They satisfy the first law
$\delta Q_{(\mathring{t})}\,\hat{=}\,\frac{\mathring{\kappa}}{8\pi G}\delta A+\mathring{\Omega}\,\delta J_{H}$ (1)
that, as we will see, can be extended to dynamical BHs.
Dynamical situations:   For dynamical BHs, there are no KHs, and EHs are unsuitable because their growth is teleological, unrelated to local physical processes (See Fig.1). But we do have quasi-local horizons (QLHs) that are free from these limitations.
[…]
Figure 1: A double Vaidya null fluid collapse. The EH $\mathcal{E}$ forms and grows in the flat region of space-time. The QLH (in green) lies entirely in the curved region. It starts out as a (space-like) DHS that grows in area in response to infalling matter, then settles down to a (null) IHS in the space-time region $M_{1}$, only to become a DHS because of the second infall and finally settles to an IHS in $M_{2}$, the only segment that coincides with the EH. Our thermodynamic considerations apply to all segments of the QLH.
Next, let us consider dynamical phases of BHs. An open portion $\mathcal{H}$ of a QLH is said to be a Dynamical Horizon Segment (DHS) if:  (i) It is nowhere null; (ii) the expansion $\theta_{(\underline{k})}$ of the other null normal $\underline{k}^{a}$ to MTSs is no where vanishing; and, (iii) a genericity condition holds: $\mathcal{E}\,\hat{=}\,|\sigma|^{2}+R_{ab}k^{a}k^{b}$ does not vanish identically on any MTS, where $\sigma_{ab}$ is the shear of $k^{a}$ and $R_{ab}$ the space-time Ricci tensor. (This condition removes some highly symmetric, degenerate cases [42, 21].) As Fig. 1 illustrates, QLHs of dynamical BHs can have multiple segments: A DHS followed by an IHS, followed by a DHS, etc. DHSs occur when there is a flux of energy into the QLH [16, 21]. IHSs occur when there is no influx. If the cosmological constant is non-negative, MTSs of DHSs are generically 2-spheres [16]. Therefore, for simplicity we will assume that the QLHs under consideration are topologically $S^{2}\times R$.
Overcoming Conceptual Obstacles:   BHs in general relativity turn out to be both simpler and more complicated than the familiar thermodynamical systems. In standard thermodynamics, a major obstacle in the passage to non-equilibrium situations is that one cannot unambiguously assign intensive parameters such as the temperature or pressure to non-equilibrium states. As we now show, this obstacle can be removed for dynamical BHs in general relativity.
Recall first that the equilibrium state $\mathfrak{e}$ of a black hole is characterized by two numbers $(R,J_{H})$. Thus the space $\mathcal{E}$ of equilibrium states is 2-dimensional. By contrast, the space $\mathcal{N}$ of non-equilibrium states is infinite dimensional since a non-equilibrium state $\mathfrak{n}$ is characterized by the plethora of fields on MTSs $S$ of a DHS $\mathcal{H}$ (that evolve from one MTS to another). One of them, the 2-metric $\tilde{q}_{ab}$ on any $S$, determines its areal radius $R[S]$ of $S$. To obtain the analog of $J_{H}$ we will restrict ourselves to DHS $\mathcal{H}$ that (are either axisymmetric or) approach a Kerr IHS in the asymptotic future (or past). Then, each $S$ also carries an infinite set of multipoles that characterize its shape and spin structure [43, 21, 31]. Among them is the angular momentum dipole $\mathcal{J}_{\mathcal{H}}[S]\,\hat{=}\,-\frac{1}{8\pi G}\oint_{S}K_{ab}\varphi^{a}dS^{b}$ where $K_{ab}$ is the extrinsic curvature of $\mathcal{H}$, and $\varphi^{a}$ an invariantly constructed rotational vector field on $\mathcal{H}$, which is: (i) tangential to and divergence-free on every MTS $S$; and, (ii) coincides with the axial Killing field if the intrinsic metric $q_{ab}$ of $\mathcal{H}$ is axisymmetric. Thus, on the space $\mathcal{N}$ of non-equilibrium states we have again a pair of observables, $(R[S],\mathcal{J}_{\mathcal{H}}[S])$, that have the same physical interpretation as $(R,J_{H})$ have on $\mathcal{E}$. Interestingly, $\varphi^{a}$ can be naturally extended off $\mathcal{H}$ so that its Komar integral $Q_{(\varphi)}[S]$ equals $2\mathcal{J}_{\mathcal{H}}$, mirroring the relation $Q_{(\mathring{\varphi})}\,\hat{=}\,-2J_{H}$ on KHs.
Therefore, there is a natural projection map $\Pi:\mathcal{N}\rightarrow\mathcal{E}$ such that $\Pi(\mathfrak{n})=\mathfrak{e}$ iff $(R,J_{H})=(R[S],\mathcal{J}_{\mathcal{H}}[S])$. Via pull-back, we can assign to any $\mathfrak{n}$, the intensive parameters $\mathring{\kappa},\,\mathring{\Omega}$ carried by $\mathfrak{e}$ (whose values are determined by $(R,J_{H})$). Thus, thanks to the BH uniqueness theorems, a key obstacle to extending thermodynamics to non-equilibrium situations is overcome: each non-equilibrium state $\mathfrak{n}$ has been assigned intensive parameters $\mathring{\kappa},\mathring{\Omega}$. Of course, $\mathring{\kappa},\,\mathring{\Omega}$ change in time on $\mathcal{H}$ since $(R[S],\mathcal{J}_{\mathcal{H}}[S])$ change from one MTS to another along $\mathcal{H}$, reflecting the fact that $\mathcal{H}$ represents a BH that is not in equilibrium.
However, as remarked earlier, BHs are also more complicated! In thermodynamics, the notion of total energy of a system in a non-equilibrium state is well-defined. In general relativity, on the other hand, the notion of energy is tied to causal vector fields. On KHs, it was natural to use the horizon Killing field $\mathring{k}^{a}$ to define energy/charge $Q_{{}^{(\mathring{k})}}$ that appears on the left side of (1). What would be its analog on DHSs? Recall that DHSs carry a preferred null direction field $k^{a}$ (with $\theta_{(k)}\,\hat{=}\,0$) which, furthermore, aligns with $\mathring{k}^{a}$ as a DHS reaches equilibrium, becoming an IHS (as in Fig. 1). Hence a natural replacement of $\mathring{k}^{a}$ would be an appropriately scaled vector $\xi^{a}$ that is parallel to $k^{a}$. Now, it was shown in [16] that, thanks to the constraint equations on the DHS, there is an invariantly defined charge $Q_{(\xi)}[S]$ associated with any such $\xi^{a}$. Suppose the infalling flux ends at a MTS $\mathring{S}$ of the QLH, so that the DHS transitions to an IHS at $\mathring{S}$. Then, the DHS and IHS assignments of charges at $\mathring{S}$ would agree iff $\xi^{a}$ is chosen such that the DHS charge $Q_{(\xi)}[\mathring{S}]$ equals the IHS charge $Q_{{}^{(\mathring{k})}}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G}|_{\mathring{S}}$ for any MTS $\mathring{S}$. Surprisingly, this consistency condition selects a unique $\xi^{a}$ on the DHS!
For this $\xi^{a}$, then, the value of the charge $Q_{(\xi)}[S]$ on any MTS $S$ of $\mathcal{H}$, equals the value of the charge $Q_{{}^{(\mathring{k})}}$ evaluated at the equilibrium state $\mathfrak{e}$ that the projection map $\Pi$ assigns to $S$. By the very definition of $\Pi$,  the agreement also holds between $\mathcal{J}_{\mathcal{H}}[S]$ and $J_{H}|_{\mathfrak{e}}$. Thus, $\Pi$ assigns to each DHS $\mathcal{H}$, a trajectory $\mathfrak{e}(R)$ in the space $\mathcal{E}$ of equilibrium states such that values of charges $(Q_{(\xi)}[S],\,\mathcal{J}_{\mathcal{H}}[S])$ on $\mathcal{H}$ agree with the values of $(Q_{{}^{(\mathring{k})}},\,J_{H})|_{\mathfrak{e}}$ all along $\mathfrak{e}(R)$! This is a surprising synergy between non-equilibrium and equilibrium values of key observables.
1st and 2nd laws on DHSs:   For definiteness, let us first focus on space-like DHSs. Then, $\hat{\tau}^{a}$ the unit normal to a DHS $\mathcal{H}$ is time-like, and $\hat{r}^{a}$ the unit normal within $\mathcal{H}$ to MTSs $S$ is space-like, and $k^{a}\,\hat{=}\,\frac{1}{\sqrt{2}}(\hat{\tau}^{a}+\hat{r}^{a})$ and $\underline{k}^{a}\,\hat{=}\,\frac{1}{\sqrt{2}}\,(\hat{\tau}^{a}-\hat{r}^{a})$ are the two null normals to $S$. As in the BCH analysis, we will use the 4 constraint equations, but now on the DHS $\mathcal{H}$ rather than on a partial Cauchy surface extending from the horizon to spatial infinity. Specifically, we will use combinations of the type $(G_{ab}-8\pi G\,T_{ab})\,\xi^{a}\hat{\tau}^{b}\,\hat{=}\,0$  of constraints, with $\xi^{a}\,\hat{=}\,\sqrt{2}\,|DR|f(R)\,k^{a}$ choosing functions $f(R)$ appropriately. (For motivation, see [16].)
On the DHS $\mathcal{H}$, $\mathcal{J}_{\mathcal{H}}[S]$ is a function of $R$, whence $\mathring{\kappa}$ and $\mathring{\Omega}$ are also functions of $R$, reflecting their dynamical nature. The unique choice of $\xi^{a}$ for which (the $\xi$-energy charge) $Q_{(\xi)}[S]\,\hat{=}\,Q_{(\mathring{k})}\equiv\frac{\mathring{\kappa}\,A}{4\pi G}$, as desired, is given by setting $f(R)\,\hat{=}\,\frac{{\rm d}}{{\rm d}R}\,(\frac{\mathring{\kappa}\,A}{2\pi G})$. Our constraint equation provides a balance law for this $Q_{(\xi)}$. Setting $\Delta\,[Q_{(\xi)}]\,\hat{=}\,Q_{(\xi)}[S_{1}]-Q_{(\xi)}[S_{2}]$, one has
$\Delta\,[Q_{(\xi)}]\,\hat{=}\,\mathfrak{F}_{(\xi)}[\mathcal{H}_{1}^{2}]\,\equiv\,\int_{\mathcal{H}_{1}^{2}}\big(\mathfrak{f}_{(\xi)}^{\,\,\rm matt}+\mathfrak{f}_{(\xi)}^{\,\,\rm gws}\big){\rm d}^{3}V.$ (2)
Here $\mathcal{H}_{1}^{2}$ is the portion of the DHS bounded by MTSs $S_{1}$ and $S_{2}$, and $\mathfrak{F}_{(\xi)}$ the flux of $\xi$-energy, with flux densities
$\mathfrak{f}_{(\xi)}^{\,\,\rm matt}\,\hat{=}\,T_{ab}\,\xi^{a}\hat{\tau}^{b}$   and,
$\mathfrak{f}_{(\xi)}^{\,\,\rm gws}\,\hat{=}\,\frac{1}{8\pi G}\,|DR|\,f(R)\,(|\sigma|^{2}+2|\zeta|^{2})$,
where $\sigma_{ab}$ is the shear of the null normal $k^{a}$, and $\zeta^{a}\,\hat{=}\,\,\tilde{q}^{ab}\,\hat{r}^{c}\nabla_{c}k_{b}$. The appearance of $|\zeta|^{2}$ seems surprising at first because it is absent in perturbative expression of the energy carried by gravitational waves across KHs. It arises because $\mathcal{H}$ is space-like rather than null: While there are only 2 true phase space degrees of freedom on null surfaces in general relativity, on a space-like surface there are 4, now captured in the fields $\sigma_{ab},\,\zeta^{a}$ tangential to MTSs, each with 2 free components [21]. They all contribute to the energy flux, as is usual in field theories.
To obtain the dynamical version of the first law (1), let us set $t^{a}:=\xi^{a}-\mathring{\Omega}\varphi^{a}$ on $\mathcal{H}$ (mimicking $\mathring{t}^{a}=\mathring{k}^{a}-\mathring{\Omega}\mathring{\varphi}$ on KHs). Then, the balance law (2) is recast as:
$\Delta\,[Q_{(\xi)}]\,\equiv\,\Delta\,[\frac{\mathring{\kappa}\,A}{4\pi G}]\,\,\hat{=}\,\,\mathfrak{F}_{(t)}[\mathcal{H}_{1}^{2}]+\mathfrak{F}_{(\mathring{\Omega}\varphi)}[\mathcal{H}_{1}^{2}].$ (3)
Matter contributions to the fluxes on the right side are obtained by replacing $\xi^{a}$ in $\mathfrak{f}_{(\xi)}^{\,\,\rm matt}$ by $t^{a}$ and $\mathring{\Omega}\varphi^{a}$ respectively. The gravitational contributions are given by [16]:
$\mathfrak{f}_{(\mathring{\Omega}\varphi)}^{\,\,\rm gws}\,\hat{=}\,\frac{1}{8\pi G}(K^{ab}-Kq^{ab})\,(\mathcal{L}_{(\mathring{\Omega}\,\varphi)}\,q_{ab})$,
leading to the equality $\mathfrak{F}_{(\mathring{\Omega}\varphi)}\,\hat{=}\,\Delta[Q_{(\mathring{\Omega}\varphi)}]$. Setting
$\mathfrak{f}_{(t)}^{\,\,\rm gws}\,:=\,\mathfrak{f}_{(\xi)}^{\,\,\rm gws}-\mathfrak{f}_{(\mathring{\Omega}\varphi)}^{\,\,\rm gws}$,
and $Q_{(t)}:=Q_{(\xi)}-Q_{(\mathring{\Omega}\varphi)}$, we have $\mathfrak{F}_{(t)}\,\hat{=}\,\Delta[Q_{(t)}]$. Since $Q_{(\mathring{\Omega}\varphi)}\,\hat{=}\,-2\mathcal{J}_{\mathcal{H}}$, the balance law (3) is equivalent to:
$\Delta\,[Q_{(t)}]\,\,\hat{=}\,\,\Delta\,[\frac{\mathring{\kappa}A}{4\pi G}]\,+\,2\Delta\,[\mathring{\Omega}\mathcal{J}_{\mathcal{H}}].$ (4)
This is the dynamical generalization of the first law (1) on KHs. There are some notable differences between the two. First, $\Delta$ refers to finite changes, and $\Delta Q_{(t)}$ and $\Delta(\mathring{\Omega}\mathcal{J}_{\mathcal{H}})$ equal the fluxes of $t$-energy and $(\mathring{\Omega}\varphi)$- angular momentum across $\mathcal{H}_{1}^{2}$ due to physical processes in space-time; this is an active version of the first law. Second, $\mathring{\kappa}$ and $\mathring{\Omega}$ are now ‘time-dependent’ –they change from one MTS to another– and therefore do not come out of $\Delta$. Third, the right side differs from that of the KH first law (1) by a factor of 2.
The second and third differences are intertwined. This becomes transparent in the infinitesimal form of the integral version, obtained by letting $S_{2}$ approach $S_{1}$ so that they are infinitesimally separated. Then our first law reduces to $\delta Q_{(t)}\,\,\hat{=}\,\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]\,+\,2\delta[\mathring{\Omega}\mathcal{J}_{\mathcal{H}}]$, where $\delta$ denotes infinitesimal changes at $S_{1}$, along $\mathcal{H}$. Since every DHS defines a trajectory  $\mathfrak{e}(R)$  on the space $\mathcal{E}$ of equilibrium states, $S_{1}$ corresponds to a point $\mathfrak{e}_{1}$ in $\mathcal{E}$, and by construction the infinitesimal changes $\delta Q_{(t)},\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]$ and $\delta[\mathring{\Omega}\mathcal{J}_{\mathcal{H}}]$ on $\mathcal{H}$ equal infinitesimal changes $\delta Q_{(\mathring{t})},\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]$ and $\delta[\mathring{\Omega}J_{H}]$ at $\mathfrak{e}_{1}$, along $\mathfrak{e}(R)$. Now, thanks to the Smarr relations on the KH, we have an identity on $\mathcal{E}$: $\frac{A\,\delta\mathring{\kappa}}{4\pi G}+2J_{H}\delta\mathring{\Omega}\,\hat{=}\,-\big(\frac{\mathring{\kappa}\,\delta A}{8\pi G}+\mathring{\Omega}\delta J\big)$. By pulling it back to $\mathcal{H}$ via $\Pi$, the infinitesimal version of the DHS first law can be rewritten as
$\delta Q_{(t)}\,\hat{=}\,\frac{\mathring{\kappa}}{8\pi G}\,\delta A\,+\,\mathring{\Omega}\,\delta\mathcal{J}_{\mathcal{H}}$ ,
which has exactly the same form as the first law (1) on KHs. But on the DHS the infinitesimal changes are caused by physical fluxes into $\mathcal{H}$ (near $S_{1}$); it is not a passive change from one equilibrium state to a nearby one. The rewriting procedure cannot be repeated to the finite version (4) of the first law because $\mathring{\kappa},\mathring{\Omega}$ are dynamical on $\mathcal{H}$ and therefore vary along the trajectory  $\mathfrak{e}(R)$  in $\mathcal{E}$.
This concludes our discussion of the generalization of the first law to dynamical BHs in general relativity. The second law for DHSs results from the combination $(G_{ab}-8\pi G\,T_{ab}){\underline{\xi}}^{a}\hat{\tau}^{b}\,\hat{=}\,0$, where ${\underline{\xi}}^{a}:=\sqrt{2}|DR|(4\pi R)\,k^{a}$. This constraint implies [16]:
$\frac{\Delta A}{4G}\,\hat{=}\,\int_{\mathcal{H}_{1}^{2}}\,\big[T_{ab}\underline{\xi}^{a}\hat{\tau}^{b}+\frac{R|DR|}{2G}(|\sigma|^{2}+2|\zeta|^{2})\big]\,{\rm d}^{3}V.$ (5)
If matter satisfies the dominant energy condition at $\mathcal{H}$, the right side is positive definite, whence the area increases. But in contrast to Hawking’s result for EHs [3], the statement of the second law on DHSs is quantitative: It is directly related to the influx of energy across the segment $\mathcal{H}_{1}^{2}$ of the DHS. As Fig. 1 shows, the area of EHs can increase across a segment even when the segment lies in a flat region of space-time where nothing is happening!
Discussion:   Space-times representing dynamical BHs do not admit KHs. They do admit EHs, but the growth in the area of EHs is unrelated to local physical processes. Therefore, the EH area cannot be a viable measure of the physical entropy in non-equilibrium situations. DHSs share some of the attractive properties with EHs –e.g., their area grows in classical GR, and they can merge but cannot bifurcate [22]. But DHSs are defined quasi-locally and are free from teleology. Therefore, they have been widely used to represent black hole far from equilibrium both und classical and quantum gravity. In this Letter we showed that they are also well-suited for extending BH thermodynamics to non-equilibrium situations. In particular, the expressions of our first and second law, and the fact that the time dependence of the area of MTSs of a DHS is governed by local physical processes, naturally lead one to identify the area of MTSs of a DHS with BH entropy in non-equilibrium situations. Finally, note that our analysis used structures that are directly available in the physical space-time. For example, one does not need bifurcate horizons that are often used, but do not exist in physical space-times of dynamical BHs beyond perturbation theory.
We will conclude with a few remarks.
1. For simplicity we focused on space-like DHSs –the most common occurrence in the classical theory [21]. The analysis goes through also for time-like DHSs [16] that e.g., represent evaporating BHs during their long semi-classical phase [33, 35, 9]. But some of the signs in various expressions flip –e.g. because of the negative energy flux into $\mathcal{H}$ during evaporation– causing the area to decrease.
2. In our discussion of the first and second law, black holes that are not in equilibrium were represented by DHSs. However, generically, in each of its dynamical phases, the BH can admit a number of ‘interweaving’ DHSs [22]. A key point is that each DHS provides us with a detailed, self-consistent description of the evolution of the BH and all our thermodynamical considerations hold for each of them. When the BH reaches equilibrium, represented by an IH, these DHSs asymptote to that IH and their thermodynamic parameters agree with those of the IH. Thus, there is consistency between the dynamical and equilibrium phases.
3. Recently, phase space methods were used to show that, already when first order dynamical perturbations are included, the entropy has to be associated not with the area of the EH, but with areas of a family of MTSs that lie inside the EH [44, 45]. The world-tube of these MTSs can be thought of as a perturbative DHS and qualitative features of the perturbative analysis can be read off from our non-perturbative expressions. However, conceptually the two frameworks are rather different. Similarities and differences will be discussed in [46]. We will also discuss charges associated with gauge fields and time-like DHSs there. However, so far our analysis has been focused on QLHs in general relativity. QLHs do exist in other gravitational theories and some of their key properties follow simply from the geometric Gauss-Codazzi equations [21]. It would be interesting to investigate whether ideas underlying our analysis can be extended to other metric theories of gravity.
Acknowledgments: This work was supported in part by the Atherton and Eberly funds of Penn State. D.E.P. acknowledges support via Penn State’s Bunton-Waller Award and the Walker fellowship from the Applied Research Laboratory at Penn State.